# Writing – review & editing, Writing – original draft, Visualization, Data curation ,Investigation, Formal analysis, Validation, Software, Methodology, Conceptualization

### v1

@@ 5.4 Attack’s Quality
Table 4 presents the results in terms of the proposed concrete metrics that utilize SHAP values to assess the quality of adversarial samples. For each attack and target model, the table shows the percentage of samples with at least one anomalous feature based on SHAP values (importance-based anomaly detection rate) and the average number of anomalous features per sample (as described in Section 4.9). Lower values indicate that the target model’s decision-making process was less affected (i.e., the model relies on similar features for prediction).
The boundary attack demonstrated high importance-based anomaly detection rates across most datasets and models, particularly on the Hate dataset, where it reached 100% across all target models except XGB (45.1%). This suggests that the Boundary attack tends to generate adversarial samples with significant changes in feature importance. The HopSkipJump attack showed variable performance, with high importance-based anomaly detection rates (e.g., 100% for the GB model on the Hate dataset) but lower rates on models such as the RF model on the ICU dataset (9.8%) and the GB model on the VideoTQ dataset (11.8%). This variability indicates that the attack’s influence on feature importance is inconsistent across models and datasets.
In general, transferability-based attacks tend to result in lower importance-based anomaly detection rates and fewer anomalous features compared to query-based attacks, indicating less distinguishable changes to feature importance. An exception to this is the random attack on the Hate and ICU datasets (100% and 80% respectively), as well as the GB importance- and LGB importance-based selection attacks (abbreviated as GB imp. and LGB imp.) on the Hate dataset, which resulted in higher importance-based anomaly detection rates (e.g., 95.1% for LGB imp. attack). However, the number of anomalous features in these cases remained relatively low, which can be explained by the fewer perturbed features, as indicated by the $L_{0}$ results ( see Section 5.2). On the ICU dataset, transferability-based attacks also resulted in lower importance-based anomaly detection rates compared to query-based attacks, with the exception of GB imp. (71.4% importance-based anomaly detection rate). This suggests that transferability-based attacks might be less disruptive in medical datasets.
The VideoTQ dataset presented mixed results, with relatively high importance-based anomaly detection rate for the XGB imp. and RF imp. attacks (up to 93.4% for XGB imp. and 100% for RF imp.), indicating variability in how transferability-based attacks affect SHAP values in communication-based datasets.
When computing the average number of anomalous features per sample, query-based attacks tend to result in a higher number of anomalous features, impacting more features across all datasets. Transferability-based attacks generally led to fewer anomalous features per sample, with the exception of the random attack, emphasizing their subtler impact on the model’s internal decision-making process.
Based on the results in Table 4, we conclude that transferability-based attacks generally produce smaller differences between the SHAP representations for benign and adversarial samples, indicating higher-quality samples with fewer anomalies and consequently, less impact on the target model’s decision-making process. However, in some instances, these anomalies are still detectable. In contrast, query-based attacks, particularly the examined Boundary and HopSkipJump attacks, significantly alter SHAP values, resulting in higher importance-based anomaly detection rates and a larger number of affected features, leading to a more pronounced impact on model behavior. Overall, SHAP-based anomaly detection proves to be an effective method for identifying adversarial samples. These findings highlight the importance of considering both the number of anomalous features and the anomaly detection rates – based on both sample values and feature importance – when assessing the impact of adversarial attacks on model robustness.
@@ 6 Discussion
The results from our evaluation of the attacker’s risk highlight the trade-offs between different attacks’ strategies. The examined query-based attacks, which are boundary and HopSkipJump attacks, achieved near-perfect success rates, often exceeding $98\%$ across all target models. However, this high consistency comes at the cost of larger distortion that is reflected in increased $L_{0}$ and $L_{2}$ distances. They are also computationally expensive due to the high volume of queries required. In contrast, the examined transferability-based attacks modified fewer features and resulted in subtle distortions, making them more efficient in terms of query overhead but less effective in terms of success rate, especially on more challenging datasets like VideoTQ. The lower attack success rates for transferability-based attacks on the evaluated datasets highlight their limitations in adversarial scenarios.
In terms of attack quality, transferability-based attacks typically produce more coherent adversarial samples than the examined query-based attacks, as indicated by lower anomaly detection rates and subtler impacts on the target model’s decision-making process, particularly on the Hate and ICU datasets. This is further supported by SHAP analysis, which reveals fewer anomalous features in these samples. These findings suggest that transferability-based attacks are inherently less detectable. The random-based selection attack is an exception, as it compromises coherence by arbitrarily modifying features, often resulting in anomalous values and noticeable shifts in SHAP values.
@@
From a defender’s perspective, however, transferability-based attacks present a unique challenge. Despite the lower transfer success rate, When adversarial samples generated by a surrogate model do manage to deceive the target model, they may be more challenging to detect, underscoring the complexity of defending against these types of attacks.
@@
Interestingly, while low $L_{2}$ distances are commonly regarded in the literature as indicative of imperceptibility, our findings suggest this may not hold for tabular data. In some cases, transferability-based attacks that achieved very low $L_{2}$ distances were still detected as anomalies by the AE. This finding calls into question the assumption that lower $L_{2}$ norms correlate with indistinguishability in tabular data domains, where imperceptibility may depend on other factors beyond the $L_{2}$ distance.
In this paper, we introduced two methods for evaluating the anomaly detection rate of black-box adversarial attacks: one based on an AE and the other on an IF model. Our results indicated that the AE model tended to identify a greater number of adversarial samples as anomalies than the IF model. This can be attributed to the nature of these models; the AE is a neural network capable of capturing more complex patterns, while the IF, a tree-based model, learns simpler patterns. These findings suggest that AE is more adept at learning benign data patterns, enabling it to identify a larger portion of adversarial samples as anomalous.
@@ 7 Conclusion
In this study we compare different attacks’ strategies of black-box adversarial attacks on ML models for tabular data. Our evaluation framework considers the attacker’s risk, effort, and the quality of adversarial samples, providing insights for both attack optimization and defense strategies.
Results across diverse target models and datasets reveal distinct trade-offs: query-based attacks, while highly effective, often sacrifice coherence by leading to significant distortions, and require substantial computational resources. Transferability-based attacks, though less successful in some cases, achieve greater coherence with fewer feature modifications and a subtler impact on model decision-making, as reflected in lower anomaly detection rates and minimal disruptions to feature importance.
Balancing coherence and effectiveness is thus essential in adversarial attack design, especially when deploying query-based attacks on models that use them as a defense strategy for tabular data. The metrics introduced in this paper to evaluate adversarial samples and attacks in the tabular domain provide a foundation for further research. This can help develop a more balanced approach that maximizes the attack effectiveness of transferability-based attacks while maintaining the input integrity of query-based attacks, ultimately aiding defenders in strengthening their defenses.
Future work may explore adversarial robustness in more complex ML pipelines in the tabular domain, such as those where models incorporate defenses against such attacks, or where the attacker does not have knowledge of all the target model’s features.

### v2

@@ 5.4 Attack Quality
Table 4 presents the results of our evaluation of attack quality using SHAP-based metrics. For each attack and target model, we report the Importance-Based Anomaly Detection Rate (percentage of samples with at least one anomalous feature based on SHAP values) and the Average Anomalous SHAP Features per Sample (as described in Section 4.9). Lower values indicate that there is less impact on the model’s decision-making process, meaning that adversarial samples were more consistent with benign samples in terms of feature importance (i.e., the model relies on similar features for prediction).
On the Hate dataset, transferability-based attacks using XGB and RF feature importance techniques (referred to in the table as XGB imp. and RF imp.) resulted in lower importance-based anomaly detection rates and lower values of average anomalous SHAP features per sample compared to other transferability- and query-based attacks. The boundary attack on the XGB model and HopSkipJump attack on the RF model also resulted in both a low importance-based anomaly detection rate and a low average number of anomalous features per sample. This indicates that the changes to feature importance were less distinguishable for those attacks.
On the ICU dataset, both query- and transferability-based attacks generally produced less anomalous SHAP values for most target models (LGB, XGB, and RF), with importance-based anomaly detection rates typically below 7% and very low values of average anomalous SHAP features per sample ranging from 0.0 to 0.1. However, when targeting the GB model, transferability-based attacks resulted in notably lower importance-based anomaly detection rates (1.8% to 10.8%) compared to query-based attacks (29.0% and 19.4% for the boundary and HopSkipJump attacks, respectively). In this case, the average number of anomalous SHAP features per sample was slightly higher for query-based attacks (0.3 to 0.4) than for transferability-based attacks (0.0 to 0.1), though still relatively low overall.
On the VideoTQ dataset, transferability-based attacks resulted in less anomalous SHAP values than query-based attacks in most cases, specifically for the GB imp., LGB imp., and RF imp. feature importance techniques, where most transferability attacks resulted in importance-based anomaly detection rates below 1%. Notably, for the GB target model, both query-based attacks resulted in lower anomalous SHAP values (5.3% and 9.4% for boundary and HopSkipJump, respectively), while some transferability-based attacks resulted in higher importance-based anomaly detection rates (51.4%, 90.3%, and 64.3% for GB imp., LGB imp., and RF imp., respectively). All attacks resulted in very few anomalous SHAP features per sample, and in most cases, the number of anomalous features per sample was close to 0.0.
Based on the results in the table, we conclude that while the examined transferability-based attacks were designed to craft subtle and well-suited perturbations and generally result in a low importance-based anomaly detection rate, they do not always outperform query-based methods in generating coherent and consistent adversarial samples. Their effectiveness depends on several key factors, including the importance-based selection method used (i.e., the model from which feature importance is computed), dataset-specific constraints and characteristics, and the target model’s architecture.
In highly constrained datasets like the ICU dataset where many features cannot be modified and over 30% of the features are categorical, transferability attacks tend to produce perturbations of similar magnitude to query-based attacks on categorical features, making query-based approaches competitive in these scenarios. The limited number of editable features forces both query- and transferability-based attacks to make fewer changes, causing less anomalous behavior overall. The random-based selection attack is an exception, as it compromises coherence by arbitrarily modifying features, often resulting in anomalous values and noticeable changes in SHAP values.
These findings suggest that defense mechanisms should consider not only the attack strategy but also the specific dataset and model vulnerabilities.
Our evaluation of attack quality—employing both the anomaly detection rate to assess sample coherence and the importance-based anomaly detection rate to evaluate the impact of adversarial samples on the target model’s decision-making process—reveals complementary insights. The Importance-Based Anomaly Detection Rate (computed based on SHAP values, see Table 4), can help identify adversarial samples that appear normal when evaluating just their feature values. For example, on the Hate dataset, the HopSkipJump attack only produced anomalous values when performed against the GB and LGB target models (see Fig. 3). However, when evaluating anomalies in SHAP values, our proposed metric captured anomalies when attacking the XGB and RF target models as well (see Table 4. On the VideoTQ dataset, we observed the opposite phenomenon; here, traditional anomaly detection (specifically using the AE model) captured anomalies for all of the transferabilty-based attacks (see Fig. 3), while the SHAP values were not anomalous in many cases (see Table 4).
These contrasting results underscore the importance of considering anomaly detection rates based on both sample values and feature importance when assessing the impact of adversarial attacks on model robustness. Each approach captures different aspects of adversarial perturbation, providing a more comprehensive understanding of the attack quality and potential defensive strategies.
5.5 Effectiveness of the Class-Specific Anomaly Detection (CSAD) Approach
Table 5: CSAD versus standard approach. The standard approach compares adversarial samples to all benign samples regardless of class, while CSAD only compares adversarial samples to benign samples of the same predicted class. Higher values indicate better detection results.
Dataset Target Model Query Attacks Transferability Attacks
boundary HopSkipJump random GB imp. LGB imp. XGB imp. RF imp.
Standard Approach (%) / Class-Specific Anomaly detection (CSAD) Approach (%)
Hate GB 87.9 / 100.0 64.1 / 96.7 94.4 / 100.0 96.7 / 100.0 100.0 / 81.2 10.0 / 27.3 32.0 / 25.6
LGB 8.6 / 58.2 40.4 / 78.0 52.6 / 64.3 50.0 / 79.7 20.0 / 76.8 22.0 / 69.9 28.6 / 40.5
XGB 0.0 / 26.4 6.6 / 67.6 70.3 / 86.4 56.0 / 92.1 50.0 / 80.3 4.5 / 22.0 13.6 / 24.8
RF 37.9 / 63.2 23.5 / 30.8 62.6 / 100.0 74.0 / 100.0 75.0 / 75.7 37.8 / 67.3 20.0 / 50.0
ICU GB 8.4 / 29.0 5.1 / 19.4 0.3 / 7.3 1.2 / 10.8 0.6 / 1.8 0.9 / 3.5 1.2 / 3.5
LGB 1.0 / 4.1 1.2 / 3.9 3.8 / 22.3 4.1 / 17.5 0.9 / 2.4 4.3 / 6.9 4.5 / 5.6
XGB 2.6 / 3.5 3.9 / 6.0 1.0 / 6.2 2.2 / 6.4 0.0 / 4.2 4.2 / 5.1 3.3 / 3.3
RF 0.0 / 0.1 0.0 / 0.2 0.0 / 4.0 0.0 / 0.5 0.0 / 0.0 0.0 / 0.0 0.0 / 0.7
VideoTQ GB 1.2 / 5.3 0.7 / 9.4 0.0 / 0.4 0.8 / 1.0 43.8 / 51.4 90.2 / 90.3 64.3 / 64.3
LGB 6.7 / 7.3 0.0 / 1.5 4.1 / 3.4 10.6 / 3.4 0.8 / 0.8 1.0 / 0.6 0.8 / 1.0
XGB 47.0 / 38.9 21.4 / 25.2 0.0 / 0.4 0.9 / 1.0 21.9 / 21.6 30.9 / 31.5 17.2 / 17.8
RF 0.0 / 66.6 0.0 / 87.6 13.8 / 0.0 17.0 / 0.0 0.0 / 0.0 0.0 / 68.8 0.0 / 79.0
Table 5 presents the results of two approaches for computing the anomaly detection rate: (1) the CSAD approach, which measures the anomalousness of SHAP values in adversarial samples in relation to benign SHAP values of the same predicted class only, and (2) the standard approach, which compares adversarial samples to the entire benign SHAP value distribution across all classes, enabling a comparative analysis of detection effectiveness between class-specific and general anomaly detection methods.
For each attack and target model, the table presents the percentage of samples with at least one anomalous feature based on SHAP values using both approaches. Higher values indicate greater success in detecting compromised adversarial samples. For each pair of results, the approach with the better detection rate is highlighted in bold. The results demonstrate that the CSAD approach substantially improves the detection of adversarial samples in most cases, often achieving a detection rate that is over twice that of the standard approach. Particularly notable are cases involving the RF target model on the VideoTQ dataset, where the CSAD approach achieved detection rates of 66.6% and 87.6% for the boundary and HopSkipJump attacks respectively, compared to 0% detection with the standard approach.
These findings confirm that assessing anomalies in data samples with respect to their specific target class is essential, particularly when the data distribution across features is not uniform. By comparing adversarial samples to benign samples of the same class, the CSAD approach accounts for class-specific patterns, leading to improved anomaly detection. This approach is especially valuable for datasets in which the normal patterns in the features of the classes differ. Subtle changes that only matter for specific classes may go undetected when a global threshold for all classes is used.
@@ 6 Discussion
The results from our evaluation of the attacker’s risk highlight the trade-offs between different attack strategies. The examined query-based attacks achieved near-perfect success rates (often exceeding $98\%$). However, this consistency comes at the cost of larger distortions and high computational overhead that is reflected in increased $L_{0}$ and $L_{2}$ distances. Transferability-based approaches offered efficiency advantages but struggled with success rates, particularly on complex datasets like the ICU dataset. The lower attack success rates for transferability-based attacks on the evaluated datasets highlight their limitations in adversarial scenarios. Importantly, when defenders rely on query-based attacks as part of defense strategies (e.g., adversarial training), it is crucial to recognize that such attacks may not capture the same degree of feature coherence exhibited by transferability-based attacks, potentially leaving gaps in robustness.
@@
On datasets with a limited number of editable features like ICU and VideoTQ, transferability attacks struggled more in terms of success rate than the more flexible Hate dataset which has a high number of editable features. However, from a defender’s perspective, transferability-based attacks present a unique challenge. Despite their lower overall success rate, when adversarial samples generated by these attacks do succeed in deceiving the target model, they may evade traditional detection mechanisms, underscoring the complexity of defending against these types of attacks.
@@
Interestingly, while small $L_{2}$ distances are commonly regarded in the literature as indicative of imperceptibility, our findings suggest this may not hold for tabular data. In some cases, transferability-based attacks that achieved very small $L_{2}$ distances were still detected as anomalies by the AE. This finding calls into question the assumption that lower $L_{2}$ norm correlate with indistinguishability in tabular data domains, where imperceptibility may depend on other factors besides the $L_{2}$ distance.
Our comprehensive evaluation, which assessed both feature space coherency and the stability of model interpretation (via SHAP), showed that adversarial perturbations can affect feature values and model interpretations independently. Some attacks (like the HopSkipJump attack when targeting XGB and RF models on the Hate dataset) preserved normal feature distributions in the sample values while significantly altering the model’s decision-making as reflected in the anomalous SHAP values. Others (like several transferability-based attacks on the VideoTQ dataset) showed the opposite pattern with anomalous sample values but relatively normal SHAP distributions. This finding suggests that comprehensive defense systems should monitor both the raw feature values and their impact on model decision-making, since focusing solely on one could cause sophisticated attacks targeting the other to go undetected.
The superiority of the CSAD approach seen in our evaluation demonstrates that context-specific thresholds substantially outperform global thresholds for identifying adversarial samples. This finding likely extends beyond our specific implementation to other defense mechanisms, where class-conditional modeling could improve detection accuracy.
Regarding anomaly detection models, the neural network-based AE detector identified more adversarial samples as anomalous than the tree-based IF alternative (as expected). These empirical results validate that AEs better capture the complex patterns and relationships in the data, making them more sensitive to subtle adversarial manipulations in tabular datasets.
@@ 7 Conclusion
In this study we compared effectiveness of different black-box adversarial attacks strategies on ML models for tabular data. Our evaluation framework considers the attacker’s risk, effort, and the quality of adversarial samples, providing insights for both attack optimization and defense strategies.
The results obtained using diverse target models and datasets highlight distinct trade-offs: query-based attacks, while highly effective, often sacrifice coherence by causing significant distortions, and require substantial computational resources. Transferability-based attacks, though less successful in some cases, achieve greater coherence and consistency with fewer feature modifications, resulting in subtler impact on model decision-making. This is evidenced by lower anomaly detection rates and minimal changes in feature importance.
Designing adversarial attacks for tabular data thus requires balancing sample coherence and the attack success rate. This balance is especially important in the context of query-based attacks, particularly when these attacks are used by defenders as part of adversarial training. To support the necessary balancing, we introduced novel metrics for assessing those aspects of adversarial samples and attacks in the tabular domain. These metrics independently capture anomalies in both feature space values and model interpretation (as reflected in SHAP values), representing the surface-level coherence of samples as well as their internal impact on the model’s decision-making process. Our evaluation framework, which combines traditional anomaly detection with SHAP-based analysis, demonstrates that perturbations can independently affect each of these aspects. This underscores the importance of monitoring both aspects when designing or defending against adversarial attacks.
Furthermore, Our CSAD approach, which employs context-specific thresholds, markedly improved detection rates over approaches that utilize a global threshold, emphasizing the importance of examining class-specific patterns in identifying adversarial samples.
These findings and metrics lay the groundwork for future research, enabling the development of more balanced attack strategies that maximize effectiveness while preserving feature space coherency. This, in turn, can help defenders bolster their systems against such threats.
Future work could investigate adversarial robustness in more complex tabular ML pipelines, including scenarios where models incorporate built-in defenses against such attacks or where attackers does not have full knowledge of the target model’s features. By building on the insights and methodologies presented here, researchers can further refine both attack and defense strategies in the tabular domain.